# Quantum Architecture Search with Unsupervised Representation Learning

### v5

@@ 1 Introduction
Quantum Computing (QC) has made significant progress over the past decades. Advances in quantum hardware and new quantum algorithms have demonstrated potential advantages [1] over classical computers in various tasks, such as image processing [2], reinforcement learning [3], knowledge graph embedding [4], and network architecture search [5, 6, 7]. However, the scale of quantum computers is still limited by environmental noise, which leads to unstable performance. These noisy intermediate-scale quantum (NISQ) devices lack fault tolerance, which is not expected to be achieved in the near future [8]. The variational quantum algorithm (VQA), a hybrid quantum algorithm that utilizes quantum operations with adjustable parameters, is considered a leading strategy in the NISQ era [9]. In VQA, the parameterized quantum circuit (PQC) with trainable parameters is viewed as a general paradigm of quantum neural networks and has achieved notable success in quantum machine learning. These parameters control quantum circuit operations, adjusting the distribution of circuit output states, and are updated by a classical optimizer based on a task-specific objective function. Although VQA faces challenges such as Barren Plateaus (BP) and scalability issues, it has demonstrated the potential to improve performance across various domains, including image processing, combinatorial optimization, chemistry, and physics [10, 11, 12]. One example of a VQA is the variational quantum eigensolver (VQE) [13, 12], which approximates the ground state and offers flexibility for quantum machine learning. We are considering using VQE to evaluate the performance of certain quantum circuits.
@@
Quantum architecture search (QAS) or quantum circuit architecture search is a framework for designing quantum circuits efficiently and automatically, aiming to optimize circuit performance [7]. Various algorithms have been proposed for QAS [5, 7, 19, 20, 6]. However, most algorithms combine the search space and search algorithm, leading to inefficiency and high evaluation costs. The effectiveness of the search algorithm often depends on how well the search space is defined, embedded, and learned. Finding a suitable circuit typically requires evaluating different architectures many times. Although predictor-based QAS [20] can separate representation learning from the search algorithm, it often relies on labeling different architectures via evaluation, and the training performance depends heavily on the quantity and quality of evaluations and the embedding. In this work, we are inspired by the idea of decoupling, and we aim to conduct QAS without labeling. We seek to explore whether decoupling can embed quantum circuit architectures into a smooth latent space, benefiting predictor-free QAS algorithms.We summarise our contributions as follows:
@@ 2 Related Work
Graph data is becoming a crucial tool for understanding complex interactions between real-world entities, such as biochemical molecules [23], social networks [24], purchase networks from e-commerce platforms [25], and academic collaboration networks [26]. Graphs are typically represented as discrete data structures, making it challenging to solve downstream tasks due to large search spaces. Our work focuses on unsupervised graph representation learning, which seeks to embed graphs into low-dimensional, compact, and continuous representations without supervision while preserving the topological structure and node attributes. In this domain, approaches such as those proposed by [27, 18, 28, 29] use local random walk statistics or matrix factorization-based objectives to learn graph representations. Alternatively, methods like [30, 31] reconstruct the graph’s adjacency matrix by predicting edge existence, while others, such as [32, 33, 34], maximize the mutual information between local node representations and pooled graph representations. Additionally, [35] investigate the expressiveness of Graph Neural Networks (GNNs) in distinguishing between different graphs and introduce Graph Isomorphism Networks (GINs), which are shown to be as powerful as the Weisfeiler-Lehman test [36] for graph isomorphism. Inspired by the success of Arch2vec [15], which employs unsupervised graph representation learning for classical neural architecture search (NAS), we adopt GINs to injectively encode quantum architecture structures, as quantum circuit architectures can also be represented as DAGs.
@@
As discussed in the previous section, PQCs are essential as ansatz for various VQAs [37]. The expressive power and entangling capacity of PQCs play a crucial role in their optimization performance [38]. Poorly designed ansatz can suffer from limited expressive power or entangling capacity, making it difficult to reach the global minimum for an optimization problem. Moreover, such ansatz may be more prone to noise [39], inefficiently utilize quantum resources, or lead to barren plateaus that hinder the optimization process [40, 41]. To address these challenges, QAS has been proposed as a systematic approach to identify optimal PQCs. The goal of QAS is to efficiently and effectively search for high-performance quantum circuits tailored to specific problems, minimizing the loss functions while adhering to constraints such as hardware qubit connections, native quantum gate sets, quantum noise models, training loss landscapes, and other practical considerations. Quantum architectures share many properties with neural network architectures, such as hierarchical, directed, and acyclic structures. As a result, QAS methods have been heavily inspired by techniques from NAS. Specifically, approaches such as greedy algorithms [42, 43], evolutionary or genetic methods [44, 45], RL-based engines [46, 47], Bayesian optimization [48], and gradient-based methods [5] have all been employed to discover improved PQCs for VQAs. However, these methods require the evaluation of numerous quantum circuits, which is both time-consuming and computationally expensive. To mitigate this issue, predictor-based approaches [19, 49] have been introduced, but they also face limitations. These approaches rely on large sets of labeled circuits to train predictors with generalized capabilities and introduce additional uncertainty into the search process, necessitating the reevaluation of candidate circuits. In this work, we propose a framework aimed at further addressing these challenges.
@@ 3 QAS with Unsupervised Representation Learning
In this work, we present our method, as illustrated in Figure 1, which consists of two independent learning components: an autoencoder for circuit architecture representation learning, and a search process that includes both search and evaluation strategies. The search space is defined by the number of gates in a circuit and an operation pool comprising general gate types such as X, Y, Z, H, Rx, Ry, Rz, U3, CNOT, CY, CZ. A random generator creates a set of circuit architectures based on predefined parameters, including the number of qubits, the number of gates, and the maximum circuit depth. These architectures are then encoded into two matrices and input into the autoencoder. The autoencoder independently learns a latent distribution from the search space and produces pre-trained architecture embeddings for the search algorithms. The evaluation strategy takes the circuit architectures generated by the search algorithm and returns a performance assessment. For evaluating circuit architectures, we use the ground state of a Hamiltonian for max-cut and quantum chemistry problems, and fidelity for quantum state preparation tasks.
3.1 Circuit Encoding Scheme
We represent quantum circuits as DAGs using the circuit encoding scheme , as described in [49, 20]. Each circuit is transformed into a DAG by mapping the gates on each qubit to a sequence of nodes, with two additional nodes added to indicate the input and output of circuits. The resulting DAG is described by an adjacency matrix, as shown in Figure 1(a). The set of nodes is further characterized by a gate matrix, which shows the node features including position information.
However, the encoding scheme represents all occupied qubits as without distinguishing between the control and target positions of two-qubit gates, which limits the effectiveness of circuit representation learning and leads to confusion during circuit reconstruction. Additionally, the adjacency matrix weights do not accurately reflect the original gate connections. To address these limitations, we propose a new encoding scheme. In our method, we explicitly encode positional information for two-qubit gates, such as CNOT and CZ, by assigning to the control qubit and to the target qubit. Furthermore, we represent the number of qubits involved in an edge as the connection weights in the adjacency matrix, as shown in Figure 1(a). These modifications enhance circuit representation learning and improve the overall effectiveness of the search.
3.2 Variational Graph Isomorphism Autoencoder
3.2.1 Preliminaries
The most common graph autoencoders (GAEs) consist of an encoder and a decoder, where the encoder maps a graph into a feature space, and the decoder reconstructs the graph from those features. One prominent example is the variational graph autoencoder (VGAE), a promising framework for unsupervised graph representation learning that utilizes a graph convolutional network as its encoder and a simple inner product as its decoder [30]. In this work, however, we do not employ the common VGAE as a framework for learning latent representations. Instead, we utilize a more powerful encoder GIN [35].
@@ 3.2.2 Encoder
The encoder GIN maps the structure and node features to latent representations . An approximation of the posterior distribution is:
where . The -layer GIN generates the embedding matrix for -layer by:
Where , and is a bias with a standard normal distribution for each layer. The is a multi-layer perceptron consisting of Linear-BatchNorm-LeakyReLU layers, and transforms a directed graph into an undirected one to capture bi-directional information. In this work, we introduce a new fusion layer, a fully connected layer that aggregates feature information from all GIN layers, rather than just the last one. The mean is computed using the fully connected layer , and similarly, the standard deviation is computed via . We can then sample the latent matrix by . For all experiments, we use GIN layers, a 16-dimensional latent vector , and a GIN encoder with hidden sizes of 128. More details on the hyperparameters can be found in Appendix A.3.
@@ 3.2.3 Decoder
The decoder takes the sampled latent variables as input to reconstruct both the adjacency matrix and the gate matrix , where encodes the gate types and encodes the qubits on which the gates act. The generative process is summarized as follows:
@@ 3.2.4 Objective Function
3.3 Architecture Search Strategies
@@ 3.3.1 Reinforcement Learning (RL)
After conducting initial trials with PPO [50] and A2C [51], we adopt REINFORCE [52] as a more effective reinforcement learning algorithm for architecture search. In this approach, the environment’s state space consists of pre-trained embeddings, and the agent uses a one-cell LSTM as its policy network. The agent selects an action, corresponding to a sampled latent vector based on the distribution of the current state, and transitions to the next state based on the chosen action. The reward for max-cut and quantum chemistry tasks is defined as the ratio of energy to ground energy, with values outside the range [0, 1] clipped to 0 or 1. For the state preparation task, circuit fidelity is used as the reward. We employ an adaptive batch size, with the number of steps per training epoch determined by the average reward of the previous epoch. Additionally, we use a linear adaptive baseline, defined by the formula , where denotes the baseline, is a predefined value in the range [0,1], and is the average reward. Each run in this work involves 1000 searches.
@@ 4 Experimental Results
To demonstrate the effectiveness and generalization capability of our approach, we conduct experiments on three well-known quantum computing applications: quantum state preparation, max-cut, and quantum chemistry. For each application, we start with a simple example involving 4 qubits and then progress to a more complex example with 8 qubits. We utilize a random generator to create 100,000 circuits as the search space, and all experiments are performed on a noise-free simulator during the search process. Detailed settings are provided in Appendix A.2. We begin by evaluating the model’s pre-training performance for unsupervised representation learning (§4.1), followed by an assessment of QAS performance based on the pre-trained latent representations (§4.2).
@@ 4.1 Pre-training Performance
Compared with subfigures 2(i), 2(j), 2(k), 2(l), 2(m), and 2(k), which utilize the encoding scheme and show more loosely distributed red points, our new encoding results in a more concentrated and smoother latent representation, as demonstrated in subfigures 2(a), 2(b) and 2(c).
@@ 4.3 Evaluation on Real Quantum Hardware
To evaluate the deployability of the circuits discovered by our architecture search, we perform additional experiments using IBM’s real quantum device ibm_sherbrooke. For both the 4-node and 8-node MaxCut tasks, we selected the best-performing circuits and trained their parameters on an noise-free simulator. Once trained, we directly executed these circuits—without any further optimization—on the real device using the same parameters.

### v6

@@ 1 Introduction
Quantum Computing has made significant progress over the past decades. Advances in quantum hardware and new quantum algorithms have demonstrated potential advantages [1] over classical computers in various tasks, such as image processing [2], reinforcement learning [3], knowledge graph embedding [4], and network architecture search [5, 6, 7]. However, the scale of quantum computers is still limited by environmental noise, which leads to unstable performance. These noisy intermediate-scale quantum (NISQ) devices lack fault tolerance, which is not expected to be achieved in the near future [8]. The variational quantum algorithm (VQA), a hybrid quantum algorithm that utilizes quantum operations with adjustable parameters, is considered a leading strategy in the NISQ era [9]. In VQA, the parameterized quantum circuit (PQC) with trainable parameters is viewed as a general paradigm of quantum neural networks and has achieved notable success in quantum machine learning. These parameters control quantum circuit operations, adjusting the distribution of circuit output states, and are updated by a classical optimizer based on a task-specific objective function. Although VQA faces challenges such as Barren Plateaus (BP) and scalability issues, it has demonstrated the potential to improve performance across various domains, including image processing, combinatorial optimization, chemistry, and physics [10, 11, 12]. One example of a VQA is the variational quantum eigensolver (VQE) [13, 12], which approximates the ground state and offers flexibility for quantum machine learning. We are considering using VQE to evaluate the performance of certain quantum circuits.
@@
Quantum architecture search (QAS) or quantum circuit architecture search is a framework for designing quantum circuits efficiently and automatically, aiming to optimize circuit performance [7]. Various algorithms have been proposed for QAS [5, 7, 19, 20, 6]. However, most algorithms combine the search space and search algorithm, leading to inefficiency and high evaluation costs. The effectiveness of the search algorithm often depends on how well the search space is defined, embedded, and learned. Finding a suitable circuit typically requires evaluating different architectures many times. Although predictor-based QAS [20] can separate representation learning from the search algorithm, it often relies on labeling different architectures via evaluation, and the training performance depends heavily on the quantity and quality of evaluations and the embedding. In this work, we decouple unsupervised architecture representation learning from the search process, enabling learned representations to be reused across different downstream tasks. This design is motivated by prior observations that the structure of parameterized quantum circuits strongly influence their expressibility, entanglement properties, and ultimately the performance of variational quantum algorithms [21, 22, 23]. Inspired by the success of Arch2vec, our approach learns structure-preserving representations of PQC architectures in an unsupervised manner, encouraging circuits with similar structural properties to occupy nearby regions in latent space, which empirically correlates with performance similarity even without labels. We seek to explore whether decoupling can embed quantum circuit architectures into a smooth latent space, benefiting predictor-free QAS algorithms. We summarise our contributions as follows:
@@ 2 Related Work
Graph data is becoming a crucial tool for understanding complex interactions between real-world entities, such as biochemical molecules [26], social networks [27], purchase networks from e-commerce platforms [28], and academic collaboration networks [29]. Graphs are typically represented as discrete data structures, making it challenging to solve downstream tasks due to large search spaces. Our work focuses on unsupervised graph representation learning, which seeks to embed graphs into low-dimensional, compact, and continuous representations without supervision while preserving the topological structure and node attributes. In this domain, approaches such as those proposed by [30, 18, 31, 32] use local random walk statistics or matrix factorization-based objectives to learn graph representations. Alternatively, methods like [33, 34] reconstruct the graph’s adjacency matrix by predicting edge existence, while others, such as [35, 36, 37], maximize the mutual information between local node representations and pooled graph representations. Additionally, [38] investigate the expressiveness of Graph Neural Networks (GNNs) in distinguishing between different graphs and introduce Graph Isomorphism Networks (GINs), which are shown to be as powerful as the Weisfeiler-Lehman test [39] for graph isomorphism. Inspired by the success of Arch2vec [15], which employs unsupervised graph representation learning for classical neural architecture search (NAS), we adopt GINs to encode quantum architecture structures, as quantum circuit architectures can also be represented as DAGs.
@@
As discussed in the previous section, PQCs are essential as ansatz for various VQAs [23]. The expressive power and entangling capacity of PQCs play a crucial role in their optimization performance [21]. Poorly designed ansatz can suffer from limited expressive power or entangling capacity, making it difficult to reach the global minimum for an optimization problem. Moreover, such ansatz may be more prone to noise [40], inefficiently utilize quantum resources, or lead to barren plateaus that hinder the optimization process [41, 42]. To address these challenges, QAS has been proposed as a systematic approach to identify optimal PQCs. Recent progress in QAS can be broadly grouped into several methodological families. Reinforcement-learning–based approaches, such as QCDS, RL-QMLAS, RL-VQC, RL-QAOA [43, 44, 45, 46], and DRL-based circuit rewriting [47], use policy agents to explore discrete circuit spaces or apply rewrite rules to improve existing circuits. Differentiable or gradient-based QAS methods have been proposed by relaxing discrete gate choices into continuous mixtures [5, 48]. While effective in settings where circuit architectures themselves can be continuously parameterized and optimized via gradients, these approaches rely on relaxation-specific assumptions and gradient access at the architecture level. In contrast, our work treats PQC architectures as discrete objects that are evaluated via task-level VQA executions, without assuming differentiability with respect to architectural choices. Exploration is instead carried out through a continuous latent embedding learned in an unsupervised manner. This distinction makes differentiable QAS complementary rather than directly comparable to the search setting considered in this work.Evolutionary methods such as QAS-EA, EQAS, EQAS-PQC, QNEAT, MoG-VQE, and NACL [7, 49, 50, 51, 52, 6, 53, 54] evolve circuit structures via mutation–selection mechanisms, often incorporating multiobjective criteria or hardware-noise awareness. Other approaches include greedy search [55, 56] and Bayesian optimization [57]. However, these methods require the evaluation of numerous quantum circuits, which is both time-consuming and computationally expensive. To mitigate this issue, predictor-based approaches [19, 58] have been introduced, but they also face limitations. These approaches rely on large sets of labeled circuits to train predictors with generalized capabilities and introduce additional uncertainty into the search process, necessitating the reevaluation of candidate circuits.
While these approaches explore circuit architectures directly in discrete or relaxed spaces, they do not leverage unsupervised representation learning to decouple structure learning from performance-driven search, which is the focus of our work. In contrast, our method differs from these lines of work in that the search is not performed directly in the discrete circuit space, nor via differentiable relaxation, but instead through a latent embedding learned in an unsupervised manner. This representation greatly reduces the combinatorial search complexity while eliminating the need for labeling or predictor training, allowing Reinforcement Learning (RL) and Bayesian Optimization (BO) to operate efficiently in a continuous latent space.
@@ 3 QAS with Unsupervised Representation Learning
In this work, we present our method, as illustrated in Figure 1, which consists of two independent components: (1) unsupervised representation learning and (2) predictor-free architecture search. The search space is defined by the number of gates in a circuit and an operation pool comprising general gate types such as X, Y, Z, H, Rx, Ry, Rz, U3, CNOT, CY, CZ. A random generator creates a set of circuit architectures based on predefined parameters, including the number of qubits, the number of gates, and the maximum circuit depth. These architectures are then encoded into two matrices and input into the autoencoder. The autoencoder independently learns a latent distribution from the search space and produces pre-trained architecture embeddings for the search algorithms. The evaluation strategy takes the circuit architectures generated by the search algorithm and returns a performance assessment. For evaluating circuit architectures, we use the ground state of a Hamiltonian for max-cut and quantum chemistry problems, and fidelity for quantum state preparation tasks.
@@
3.1 Circuit Encoding Scheme
Following Definition 1, we represent quantum circuits as DAGs, extending the encoding strategy used in [49,20]. Each quantum gate in the circuit is treated as a node, and directed edges encode execution dependencies between gates acting on the same qubit. Because a quantum circuit is a sequence of quantum operations, we additionally introduce a global start node and a global end node so that all qubit lines originate from the former and terminate at the latter. This guarantees a consistent topological ordering and ensures that the resulting structure forms a DAG. From this DAG, we construct two complementary representations. The adjacency matrix captures the structural topology of the circuit, where if there exists a directed edge from node to node , and otherwise. In parallel, we construct the gate matrix , where each row corresponds to a gate and is encoded using a one-hot or embedding-based representation of its gate type.
Limitations of prior encodings.
However, the original encoding treats all occupied qubits uniformly and does not distinguish control and target roles in two-qubit operations (e.g., CNOT, CZ). This ambiguity may lead to representation loss and suboptimal circuit reconstruction. Moreover, the adjacency matrix in the original scheme assigns uniform edge weights, which fails to reflect the structural importance of multi-qubit interactions.
Enhanced encoding scheme.
To address these limitations, we introduce an enhanced encoding scheme. For two-qubit gates as shown in Figure 1(a), we explicitly encode operational roles by assigning to the control qubit and to the target qubit in the corresponding row of . Additionally, we assign edge weights proportional to the number of qubits involved in the operation (e.g., weight for two-qubit gates) in the adjacency matrix. These modifications retain operational structure, qubit roles, and gate-type distinctions, enabling a more expressive and faithful circuit representation. As demonstrated in later sections, this refined encoding improves unsupervised representation learning and enhances the performance of latent-space architecture search.
3.2 Unsupervised representation learning
The objective of this stage is to learn latent representations of quantum circuits without relying on performance labels or VQA evaluations. We jointly train a GIN-based encoder and a fully connected decoder by maximizing the ELBO over 100,000 randomly generated circuits using only structural inputs .
3.2.1 Preliminaries
Graph autoencoders (GAEs) provide a general framework for unsupervised graph representation learning, consisting of an encoder that maps a graph to a latent space and a decoder that reconstructs the graph from this representation. A representative example is the variational graph autoencoder (VGAE), which uses a graph convolutional network as the encoder and optimizes a variational objective [33].
In this work, we adopt the graph autoencoder paradigm but depart from the standard VGAE design. Specifically, we employ a Graph Isomorphism Network (GIN) [38] as the encoder, which has been shown to be theoretically more expressive than common GCN-based encoders in distinguishing graph structures. This choice is particularly important for quantum circuits, where subtle differences in connectivity and gate ordering can lead to significant differences in circuit behavior.
@@ 3.2.2 Encoder
The encoder GIN maps the circuit structure and node features to a set of latent variables . We approximate the posterior distribution over latent variables as
where . During representation learning, this variational posterior provides circuit-dependent latent encodings, while after training latent vectors are sampled from the prior for architecture search.
The GIN consists of message-passing layers, which generate intermediate node embeddings according to:
where denotes the initial node feature matrix derived from the gate encoding. The term performs neighborhood aggregation by summing feature vectors of gates that are connected in the circuit graph, thereby propagating structural information along circuit dependencies. The term introduces a learnable self-loop that controls the relative contribution of a node’s own features at layer . The symmetrized adjacency matrix enables bidirectional message passing between connected gates. Finally, applies a shared nonlinear transformation to the aggregated features to produce the updated node embeddings at layer .
After GIN layers, we aggregate the learned node embeddings to parameterize the variational posterior. Instead of relying solely on the final GIN layer, we introduce a fusion layer that aggregates information from all layers to improve representation robustness. The mean and standard deviation are computed as and , respectively. Latent variables are then sampled using the reparameterization trick , where .
For all experiments, we use GIN layers, a latent dimension of , and a hidden dimension of for the GIN encoder. Additional hyperparameter details are provided in Appendix A.4.
@@ 3.2.3 Decoder
The decoder takes the sampled latent variables as input to reconstruct both the adjacency matrix and the gate matrix , where corresponds to the one-hot encoding of the gate type for each node, and encodes the qubit positions associated with each gate. The generative process is summarized as follows:
@@ 3.2.4 Objective Function
The representation learning stage is fully unsupervised and operates solely on circuit structure . Because the GIN encoder aggregates local gate patterns and connectivity, circuits with similar structural characteristics tend to yield similar node embeddings, and the reconstruction objective encourages these embeddings to form a smooth latent space. This intuition aligns with prior studies showing that circuit structure strongly influences expressibility and entangling capability, and thereby the expressive behavior of PQCs in variational algorithms [21, 22, 23]. As observed in Arch2vec, a variational graph isomorphism autoencoder trained only on architectural structure (without using performance labels) encourages architectures with similar connections to cluster in latent space, and architectures with similar performance tend to occupy nearby regions in that space [15]. As a result, shared structural patterns often correlate with performance-relevant properties in variational algorithms, an effect that is also reflected in the smooth organization of circuits observed in our latent space visualizations (Figure 2).
3.3 Predictor-free architecture search
In the search stage, we freeze the pretrained encoder and explore the latent space using RL/BO. Candidate architectures are evaluated only for reward computation, resulting in a predictor-free search process.
@@ 3.3.1 Reinforcement Learning (RL)
After conducting initial trials with PPO [59] and A2C [60], we adopt REINFORCE [61] as a more effective reinforcement learning algorithm for architecture search. In this approach, the environment’s state space consists of pre-trained embeddings, and the agent uses a one-cell LSTM as its policy network. The agent selects an action, corresponding to a sampled latent vector based on the distribution of the current state, and transitions to the next state based on the chosen action. For the MaxCut and quantum chemistry tasks, we adopt an energy convention in which the ground-state energy is negative and corresponds to a trivial upper bound. The reward used in the REINFORCE search is defined as
Under this convention, for valid evaluations, where indicates reaching the ground state and corresponds to . During early optimization or in the presence of numerical noise, may become slightly positive, producing a negative ratio; in such cases, we clip the reward to to ensure a stable and interpretable reward signal. For the state preparation task, circuit fidelity is used as the reward. We employ an adaptive batch size, with the number of steps per training epoch determined by the average reward of the previous epoch. Additionally, we use a linear adaptive baseline, defined by the formula , where denotes the baseline, is a predefined value in the range [0,1], and is the average reward. Each run in this work involves 1000 searches.
@@ 4 Experimental Results
To demonstrate the effectiveness and generalization capability of our approach, we conduct experiments on three well-known quantum computing applications: quantum state preparation, max-cut, and quantum chemistry. For each application, we start with a simple example involving 4 qubits and then progress to a more complex example with 8 qubits. We utilize a random generator to create 100,000 circuits as the search space, and all experiments are performed on a noise-free simulator during the search process. Detailed settings are provided in Appendix A.2. We begin by evaluating the model’s pre-training performance for unsupervised representation learning (Appendix 4.1), followed by an assessment of QAS performance based on the pre-trained latent representations (Appendix 4.2).
@@ 4.1 Pre-training Performance
Compared with subfigures 2(i), 2(j), 2(k), 2(l), 2(m), and 2(n), which utilize the encoding scheme and show more loosely distributed red points, our new encoding results in a more concentrated and smoother latent representation, as demonstrated in subfigures 2(a), 2(b) and 2(c).
@@ 4.2 Quantum Architecture Search (QAS) Performance
We further study the robustness of our framework with respect to the latent dimension and the depth of the GIN encoder; ablation results for both RL- and BO-based search are reported in Appendix A.3.
@@ 4.3 Evaluation on Real Quantum Hardware
To evaluate the deployability of the circuits discovered by our architecture search, we perform additional experiments using IBM’s real quantum device ibm_sherbrooke. For both the 4-node and 8-node MaxCut tasks, we selected the best-performing circuits and trained their parameters on a noise-free simulator. Once trained, we directly executed these circuits—without any further optimization—on the real device using the same parameters.